Authentication Server Hiccup.

Tue May 4 15:25:44 PDT 2004 — Authentication Server Hiccup. Our authentication servers briefly failed all authentication attempts due to a corrupted automatic configuration update. We are tracking down the cause of the corruption so we can resolve whatever issues may have caused it. -Kelsey and Russ

Sasser and Gaobot on the loose.

Mon May 3 09:52:32 PDT 2004 — Sasser and Gaobot on the loose. Over the weekend, two new viruses have been infecting numerous hosts on the Internet. They exploit vulnerabilities in Windows NT, Windows 2000, Windows Server 2003, and Windows XP. Other systems should be unaffected. Symptoms of infection include severe degradation of computer performance and error messages regarding “isass.exe” or “lsass.exe”. Four variants of the Sasser worm have appeared since it was first detected late last week.

Removal resources can be found at the following URLs: sarc.com/avcenter/venc/data/w32.sasser.worm.html www.sarc.com/avcenter/venc/data/w32.gaobot.afw.html vil.nai.com/vil/stinger/

As always, Sonic.net highly recommends running Windows Update frequently on any Microsoft Windows systems. — John F and the tech support group

DSL maintenance complete.

Mon May 3 00:46:24 PDT 2004 — DSL maintenance complete. The scheduled circuit upgrade is complete, and all DSL customers out of our San Francisco 200 Paul POP are up on the new OC3. Actual downtime was between 10 and 30 minutes. We also used the maintenance window to upgrade the software on the San Francisco DSL concentrator. -Nathan and John

DSL maintenance outage May 3.

Thu Apr 29 12:45:57 PDT 2004 — DSL maintenance outage May 3. We are going to upgrade one of our DSL trunks from SBC/ASI and the equipment that terminates it on Monday, May 3, 2004, at 12:01 AM. Customers who will be affected are the ones served from our San Francisco 200 Paul POP. The duration of the outage will vary by customer from 10 minutes to 2 hours as ASI reroutes our circuits from the old DS3 to the new OC3. -John and Nathan

Graton Rooftop Outage.

Tue Apr 27 16:13:59 PDT 2004 — Graton Rooftop Outage. The Graton Rooftop network is offline. The problem appears to be with the head-end unit as other hosts at the tower have connectivity. We are working quickly to resolve the issue. -Matt and Bryan

Occidental Rooftop Outage.

Mon Apr 26 09:05:25 PDT 2004 — Occidental Rooftop Outage. An internal wiring issue has created a power outage in Occidental and caused our our head-end wireless router to go off-line. An electrician is on the way to fix the problem. There is no ETR at this time. -Matt

Update: Mon Apr 26 09:24:35 PDT 2004 — Power has been restored. Power has returned to the site and Rooftop service has been restored. -Matt

Night Operations Complete.

Thu Apr 22 02:52:32 PDT 2004 — Night Operations Complete. Most of the planned work was accomplished without event. sonic.sonic.net took a bit longer to get to reboot on the new kernel than expected and while we didn’t replace the gigabit switch we did shuffle around some of our layer three switches allow for expansion. -Kelsey, Nathan, Kevan and Jared

Operating system upgrades across the board.

Wed Apr 21 12:01:34 PDT 2004 — Operating system upgrades across the board. We are upgrading kernels throughout our network do to a recent discovery of a kernel vulnerability with Linux. This is a “local root” vulnerability, which means we needed to upgrade all systems with customer access before posting about this in our MOTD. This has now been done. There were brief outages in single-server services, such as our shell server, as the systems were rebooted.

Administrators of Linux systems with multiple users may well be served by reading the following link:

www.securityfocus.com/archive/1/360810/2004-04-18/2004-04-24/0

-Scott, Geoff, and (for most of our systems!) Kelsey

Night Operations.

Wed Apr 21 16:24:35 PDT 2004 — Night Operations. Tonight, starting at Midnight, we will be rebooting the remaining servers to finish upgrading the kernels. In addition to this, we and also going to be servicing our VPN concentrator and replacing one of our gigabit ethernet switches. Some services may be briefly affected while servers are rebooted or moved to the new switch. -Kelsey, Nathan, Kevan and Jared.

Outbound Email Bulk Detection.

Wed Apr 21 16:07:14 PDT 2004 — Outbound Email Bulk Detection. We enabled outbound Email bulk this afternoon and the system is functioning as expected. If you send out broadcast email traffic and wish to be excepted from the detection, please contact support@sonic.net with an explanation detailing the type of mail you send, how frequently, to how many recipients and other bits of information that you feel are relevant. -Kelsey